What is the value of x in the solution to the system y = 4x + 3 and 3x + 2y = -16? (Round to the nearest integer if necessary.)
Answer by sofiyac(983)

3x+2(4x+3)=-16
3x+8x+6=-16
11x+6=-16
11x=-22
x=-2
so the value of x is -2